1.It is 12pm.2 people are fishing.If Benoit gives 2 of his fishes to Amelie,she will have 2 times more fishes than him. If Amelie gives 4 fishes to Benoit,he will have the same amount of fishes.How many fishes does each have ??

2.It is 4pm.If benoit gives 8 of his fishes to Amelie,she will have 3 times more fishes than him.If Amelie gives the 1/5 of her fishes to Benoit,they willhave the same amount.Howmany fishes do the have?? PLESEEEEEEEEEEEEEE HELPPPPPPPPPPPPPPPPPP WITH THIS ONE !!!!!!!!!!!!!!!!!!!!!!!!! THANK YOU SOOOOOOOOOOOOOOOO MUCHHHHHHHHHHHHHHH !!!!!!!!!!!!!!!! OH GOD !!!!!!!!!!!!

Answer by stanbon(75887)

1.It is 12pm.2 people are fishing.If Benoit gives 2 of his fishes to Amelie,she will have 2 times more fishes than him.
-------------
Number of fish they each start with:
Amelia has "a"
Benoit has "b"
-------------------
After switching the fish the number of fish for each is:
Amelie has "a+2"
Benoit has "b-2"
------
"she will have 2 times more fishes than him".
---
Resulting Equation:
a+2 = 2(b-2)
a -2b = -6
-------------------------------------
If Amelie gives 4 fishes to Benoit,he will have the same amount of fishes.
Amelie: a+2-4 = a-2
Benoit: b-2+4 = b+2
Resulting Equation:
a-2 = b+2
a - b = 4
----------------
"How many fishes does each have ??
Solve the system of equations:
a - 2b = -6
a - b = 4
-----------------
Subtract 1st equation from the 2nd to get:
b = 10 (# of fish Benoit has to start with)
b+2 = 12 (# of fish Benoit has at 4 o'clock.
---
Substitute to solve for "a".
a - 10 = 4
a = 14 (# of fish Amelie has to start with)
a-2 =12 (# of fish she has at 4 o'clock.
============================================================================
2.It is 4pm.If benoit gives 8 of his fishes to Amelie,she will have 3 times more fishes than him..How many fishes do they have?
---
b = 12 -8 = 4 (# that Benoit has after giving 8 to Amelie)
a = 12 +8 = 20(# that Amelie has after receiving 8 from Benoit)
I guess they are saying a = b + 3b; 20 = 12 + 3*4.
But they have a strange way of saying it.
===============================================================
Cheers,
Stan H.